Which choice has a greater present value if we assume a required rate of return of 10%?

1: A lump sum cash flow today of $248.69

2: $100 cash flows occurring one, two, and three years from today

3: a single cash flow of $331 three years from today.

a. Choice 1

b. Choice 2

c. Choice 3

d. The choices all have equal present values at a discount rate of 10%.
Answer #1

d. The choices all have equal present values at a discount rate of 10%.

PV are as follows

Option 1 = 248.69

Option 2 = 100/1.1 + 100/1.1^2 + 100/1.1^3 = 248.69

Option 3: 331/1.1^3 = 248.69

All are the same present values.